Given: V is constant, M grows 10% per year, Y grows 4% per year, r=8.
a). Solve for ^
b). Solve for ¡
c). If the central bank increase the money growth rate by 4 percentage pants per year, dind ∆¡
d). Suppose the growth rate of Y falls 2% per year. What will happen π?
e). What must the central bank do if it wishes to keep π constant?
1
Expert's answer
2020-09-21T08:11:31-0400

a) π = 10% - 4% = 6%.

b) i = r - π = 8 - 6 = 2%.

c) If the central bank increase the money growth rate by 4 percentage points per year, then ∆i = 4%.

d) If the growth rate of Y falls 2% per year, then π will decrease too.

e) The central bank should decrease M if it wishes to keep π constant.